# Enabling Automatic Differentiation with Mollified Graph Neural Operators

### v1

@@ 2.3 Computing Derivatives
In order to use physics losses, a major technical challenge is to efficiently compute derivatives without sacrificing accuracy, since numerical errors made on the derivatives will be amplified in the physics losses and output solution.
Finite Differences (FD).
 A simple approach is to use numerical derivatives computed using finite differences (FD). This differentiation method is fast and memory-efficient: given a $n$-points grid, it requires $O(n)$ computations. However, numerical differentiation using FD faces the same challenges as the corresponding numerical solvers: it requires a fine-resolution grid to be accurate and therefore becomes intractable for multi-scale and fast-varying dynamics. On point clouds, the stencil coefficients in FD formulas vary from point to point and must be computed each time, as outlined in Appendix G, adding to the computational cost. The errors in the resulting derivatives will also vary across the domain depending on the density of nearby points.
@@
 Fourier differentiation is also fast and memory-efficient to approximate derivatives as it requires $O(n\log n)$ given a $n$-points grid. However, just like spectral solvers, it requires smoothness, uniform grids, and performs best when applied to periodic problems. Fourier differentiation can be performed on non-uniform grids but the computational cost grows to $O(n^{2})$, and if the target function is non-periodic or non-smooth, the Fourier differentiation is not accurate. To deal with this issue, the Fourier continuation method (Maust et al., 2022) can be applied to embed the problem domain into a larger periodic space, at the cost of higher computational and memory complexity.
@@
 Derivatives can be computed pointwise using automatic differentiation by applying the chain rule to the sequence of operations in the model. Autograd (Maclaurin et al., 2015) automates this process by constructing a computational graph during the forward pass and leveraging reverse-mode differentiation to compute gradients during the backward pass. Autograd is typically the preferred method for computing derivatives in PINNs for a variety of reasons (Baydin et al., 2017): (1) Unlike FD, which introduces discretization errors, autograd provides exact derivatives (up to the limits of machine numerical precision), ensuring the accuracy that is critical in physics-informed machine learning, where derivative errors are amplified in physics losses. (2) Autograd computes gradients in a single pass (while FD requires multiple function evaluations). It provides exact derivatives regardless of the mesh resolution, making it particularly advantageous for large-scale problems, where FD become computationally intractable. (3) Autograd can compute higher-order derivatives with minimal additional cost, while FD require additional function evaluations and can suffer from further error accumulation. (4) Autograd can compute derivatives at any point in the domain seamlessly, while FD can struggle to handle complex geometries. (5) Figure 2 empirically shows that autograd derivatives are smoother and more stable than FD with the proposed model. However, autograd also has limitations: all operations need to be differentiable, and storing all intermediate computations in a computational graph can significantly increase memory usage for deep models. As a result, it can be slower and more memory-intensive than FD for simpler low-dimensional problems for which FD are accurate enough at low resolutions. When the physics loss involves a deep composition of operations, issues of vanishing or exploding gradients can also be exacerbated during backward propagation with automatic differentiation.
@@ 3 Methodology
 Recall from Equation 3 that GNO computes the integral of an indicator function $\mathds{1}_{\text{B}_{r}(x)}$ that is not differentiable. We propose a fully differentiable layer that replaces the indicator function $\mathds{1}_{\text{B}_{r}(x)}$ with a differentiable weight function $w$ supported within $\mathds{1}_{\text{B}_{r}(x)}$. This is inspired by the mollifier in functional analysis (Evans, 2010), which is used to smooth out the indicator function. The resulting mollified GNO ($m$GNO) acts on an input function $v$ via
@@
which allows to recover exact derivatives at the boundary. Then, we can compute the derivative
@@
Automatic differentiation algorithms can then be used to compute the derivatives appearing in physics losses. We can also use a cached version of neighbor search (i.e. store neighbors with nonzero weight) to keep the method as efficient as the original GNO up to the negligible cost of evaluating the weight function $w$. We emphasize that these definitions work for arbitrarily complex domains $D$.
As for the $m$GNO, the radius $r$ is a hyperparameter and the integral can be approximated with a Riemann sum, for instance. An example of simplified pseudocode for the $m$GNO layer is provided in Appendix C.
@@
The mollified GNO can then be used within a fully differentiable mollified GINO ($m$GINO) to learn efficiently the solution operator of PDEs with varying geometries using physics losses where the derivatives are computed using automatic differentiation,
@@
This allows $m$GINO to be used for solving inverse design and shape optimization problems on complex geometries.
@@
 Letting $d=\|x-y\|^{2}/r^{2}$, examples of weight functions $w$ with support in $\mathds{1}_{\text{B}_{r}(x)}(y)$ are given by
@@
These weighting functions, displayed in Figure 10, are decreasing functions from $1$ to 00 on $[0,r]$.
4 Experiments
We use Meta-PDE (Qin et al., 2022) and the popular finite element method (FEM) FEniCS (Alnæs et al., 2015; Logg et al., 2011) as baselines. Meta-PDE learns initializations for PINNs over multiple instances that can be finetuned on any single instance, and two versions have been proposed based on the meta-learning algorithms MAML (Finn et al., 2017) and LEAP (Flennerhag et al., 2019). While FEniCS has been successfully applied in various disciplines, it has not been highly optimized for our applications and could possibly be outperformed by other FEMs and non-FEMs (Liu, 2009) such as Radial Basis Function (RBF) method (interpolation using RBFs), Finite Point Methods (FPMs) and Moving Least Squares (MLS) methods (weighted least squares to approximate solutions) and spectral methods (expanding the solution in a basis of functions).
@@ 4.1 Burgers’ Equation
Table 1: Relative L2 data loss and physics loss (computed using Autograd) for the proposed $m\mathrm{GNO}\circ\mathrm{FNO}$ models and PINO baselines, trained with $\mathcal{L}_{\text{Burgers}}$ where derivatives are computed using different methods.
Physics Loss Data Loss
Autograd (ours) $1.62\cdot 10^{-6}$ $\mathbf{1.33}\cdot\mathbf{10^{-2}}$
@@
We consider the 1D time-dependent Burgers’ equation with periodic boundary conditions, whose description is provided in greater detail in Section D.1. We wish to learn the mapping $\mathcal{G}^{\dagger}$ from the initial condition $u(x,0)=u_{0}$ to the solution $u(x,t)$. Examples of solutions are shown in Figure 3. The input initial condition $u_{0}(x)$ given on a regular spatial grid is first duplicated along the temporal dimension to obtain a 2D regular grid, and then passed through a 2D FNO and a mollified GNO to produce a predicted function $v=(\mathcal{G}_{m\mathrm{GNO}}\ \circ\ \mathcal{G}_{\mathrm{FNO}})(u_{0})$ approximating the solution $u(x,t)$. We minimize a weighted sum $\mathcal{L}_{\mathrm{Burgers}}$ of the PDE residual and initial condition loss (see Equation 17).
We trained models using $\mathcal{L}_{\text{Burgers}}$, where the derivatives are computed using various differentiation methods. The results, obtained by finetuning the models one instance at a time and then averaging over the dataset, are presented in Table 1. Compared to numerical differentiation, $m\mathrm{GNO}\circ\mathrm{FNO}$ with Autograd was easier to tune and produced a model with data loss 20× lower (although 6× slower per training epoch), suggesting it better captures the physics underlying the data. It also achieved a physics loss 3000× smaller, but this was expected since these were evaluated using autograd.
@@
A comparison to PINO (i.e. using a FNO instead of a $m\mathrm{GNO}\circ\mathrm{FNO})$ shows that PINO achieved a lower physics loss, but failed to reduce the Relative L2 data loss below $10\%$ on the hyperparameter sweep considered. Note that autograd $m\mathrm{GNO}\circ\mathrm{FNO}$ was only 1.2× slower than autograd PINO (although this was with a batch size of 1 and autograd PINO could be accelerated more easily using batching).
@@
Table 2 displays the balance between training with physics and/or data losses. Using the data loss helps reduce the gap with reference data, but can come at then cost of a higher physics loss, and vice-versa.
We also tried pretraining the $m\mathrm{GNO}\circ\mathrm{FNO}$ models using the data loss before finetuning with $\mathcal{L}_{\text{Burgers}}$, but were not able to obtain better results this way. The data loss rapidly deteriorates from its original value obtained in a data-driven way, while the physics loss improves slowly but does not get better than the hybrid and physics only versions.
We also considered randomly subsampling the points at which the PDE residuals are evaluated, and Table 3 shows that despite the random location and reduction of the number of points used to evaluate the derivatives, we maintain good results, until the number of points becomes very low.
@@
Table 2: Data loss (relative L2) and PDE loss of three models trained with different losses $\mathcal{L}$, using autograd to compute the physics losses, for the Burgers’ equation.
Data $1.42\cdot 10^{-3}$ $3.81\cdot 10^{-3}$ $1.33\cdot 10^{-2}$
PDE $4.80\cdot 10^{-4}$ $1.48\cdot 10^{-4}$ $1.62\cdot 10^{-6}$
@@
Table 3: Physics and data losses (relative L2) for models trained with the physics loss $\mathcal{L}_{\text{Burgers}}$ using autograd derivatives evaluated at different numbers of randomly subsampled points from the original $128\times 26$ regular grid of $3328$ points)
Physics Loss Data Loss
@@
Random 250 points $1.47\cdot 10^{-5}$ $1.44\cdot 10^{-2}$

### v2

@@ 2.3 Computing Derivatives
To use physics losses, a major technical challenge is to efficiently compute derivatives without sacrificing accuracy, since numerical errors on the derivatives will be amplified in the physics losses and output solution.
Finite Differences.
A simple approach is to use numerical derivatives computed using finite differences. This differentiation method is fast and memory-efficient (it requires $O(n)$ computations for an $n$-point grid). However, it faces the same challenges as finite difference numerical solvers: it requires a fine-resolution grid to be accurate and therefore becomes intractable for multi-scale and fast-varying dynamics. On point clouds, the stencil coefficients in finite difference formulas vary from point to point and must be computed each time, as outlined in Appendix˜E, adding to the computational cost. The errors in the resulting derivatives will also vary across the domain depending on the density of nearby points.
@@
Fourier differentiation is also fast and memory-efficient to approximate derivatives as it requires $O(n\log n)$ given an $n$-point grid. However, just like spectral solvers, it requires smoothness, uniform grids, and performs best when applied to periodic problems. Fourier differentiation can be performed on non-uniform grids, but the computational cost grows to $O(n^{2})$, and if the target function is non-periodic or non-smooth, the Fourier differentiation is not accurate. To deal with this issue, the Fourier continuation method (Ganeshram et al., 2025) can be applied to embed the problem domain into a larger periodic space, at the cost of higher computational and memory complexity.
@@
Derivatives can be computed pointwise using automatic differentiation by applying the chain rule to the sequence of operations in the model. Autograd (Maclaurin et al., 2015) automates this process by constructing a computational graph during the forward pass and leveraging reverse-mode differentiation to compute gradients during the backward pass. Autograd is typically the preferred method for computing derivatives in PINNs for a variety of reasons (Baydin et al., 2017):
  • Unlike finite differences, which can introduce large discretization errors when not computed on fine enough grids, Autograd provides exact derivatives, ensuring the accuracy that is critical in physics-informed machine learning, where derivative errors are amplified in physics losses. By exact derivatives, we mean that automatic differentiation computes gradients precisely from the differentiable computation graph, accurate up to machine precision. While the graph represents a discretized approximation of the continuous PDE, in practice, the computed loss and gradients are extremely accurate, and automatic differentiation remains highly reliable even when evaluated on a relatively small set of points, as demonstrated in our numerical experiments.
  • Autograd computes gradients in a single pass (while finite differences require multiple function evaluations). It provides exact derivatives regardless of the mesh resolution, making it particularly advantageous for large-scale problems, where finite differences become computationally intractable.
  • Autograd can compute higher-order derivatives with minimal additional cost, while finite differences require additional function evaluations and can suffer from further error accumulation.
  • Autograd can compute derivatives at any point in the domain seamlessly, while finite differences can struggle to handle complex geometries.
  • Figure˜3 empirically shows that Autograd derivatives are smoother and more stable than finite differences with the proposed model.
However, Autograd also has limitations: all operations need to be differentiable, and storing all intermediate computations in a computational graph can significantly increase memory usage for deep models. As a result, it can be slower and more memory-intensive than finite differences for simpler low-dimensional problems for which finite differences are accurate enough at low resolutions. When the physics loss involves a deep composition of operations, issues of vanishing or exploding gradients can also be exacerbated during backward propagation with automatic differentiation.
Figure 3: Qualitative assessment of the smoothness of derivatives. We fit mmGINO with a physics loss to the differentiable function $u(x,y)=\sin(4\pi xy)$ on $0,1]^{2}$, and compare the analytic ground truth derivative $\partial_{xx}u(x,y)=-16\pi^{2}y^{2}\sin(4\pi xy)$ (left) with the numerical derivatives obtained using automatic differentiation (middle) and finite differences (right).
@@ 3 Methodology
Recall from Equation˜3 that the GNO computes the integral of an indicator function $\mathds{1}_{\text{B}_{r}(x)}$ that is not differentiable. We propose a fully differentiable layer that replaces the indicator function $\mathds{1}_{\text{B}_{r}(x)}$ with a differentiable weight function $w$ supported within $\text{B}_{r}(x)$. This is inspired by the mollifier in functional analysis (Evans, 2010), which is used to smooth out the indicator function. The resulting mollified graph neural operator (mmGNO) acts on an input function $v$ via
@@
which allows the recovery of exact derivatives at the boundary. Then, we can compute the derivative
@@
Automatic differentiation algorithms can then be used to compute the derivatives appearing in physics losses. We can also use a cached version of neighbor search (i.e. store neighbors with nonzero weight) to keep the method as efficient as the original GNO up to the negligible cost of evaluating the weight function $w$. We emphasize that these definitions work for arbitrarily complex domains $D$. As for the GNO, the radius $r$ is a hyperparameter and the integral can be approximated with a Riemann sum, for instance. An example of simplified pseudocode for the mmGNO layer is provided in Appendix˜D.
@@
The mollified GNO can then be used within a fully differentiable mollified geometry-informed neural operator (mmGINO) to learn efficiently the solution operator of PDEs with varying geometries using physics losses where the derivatives are computed using automatic differentiation,
@@
This allows mmGINO to be used for solving inverse design and shape optimization problems on complex geometries, by backpropagating gradients with respect to control parameters through it.
@@
The choice of weight function plays a central role in mollifying the hard cutoff neighborhoods of standard GNOs with smooth, compactly supported functions that enable stable and accurate automatic differentiation. These weights gradually attenuate contributions near the neighborhood boundary, improving gradient stability and reducing artifacts. Key properties include smoothness to avoid spurious gradient oscillations, compact support to control computational cost by ignoring distant points, and locality tuned through the radius $r$.
In practice, smooth bell-shaped or cosine-like profiles often perform well and are well-suited. The radius $r$ is typically chosen to obtain a desired typical number of neighbors per query point which balances efficiency and information aggregation. Performance is poor when the radius is too small, and when $r$ becomes too large, the computational and memory costs significantly increase while performance deteriorates. In practice, one can start from a smaller value of $r$ and progressively increase it until satisfactory performance is achieved.
Letting $d=\|x-y\|_{2}/r$, examples of weight functions $w$ with support in $\mathds{1}_{\text{B}_{r}(x)}(y)$ are
@@
These weight functions, displayed on the right in Figure˜1, are decreasing functions from $1$ to 0 on $[0,r]$, and preserve differentiability for all points.
An ablation study for the radius $r$ and choice of weight function $w$ is conducted for the nonlinear Poisson equation in Section˜4.2.6, and a discussion on the choice of aggregation scheme for neighbors contributions is provided in Section˜4.2.3, in particular to highlight how a mean aggregation can lead to undesired artifacts when training based on a data loss only.
Remark. The entire mmGNO and mmGINO architectures are designed such that all of their components are differentiable, ensuring that derivatives of any order can be computed correctly and accurately using automatic differentiation. For instance, we avoid non-differentiable activation functions such as ReLU (Rectified Linear Unit), which introduce discontinuities in their derivatives, and instead use smooth alternatives like GeLU (Gaussian Error Linear Unit). All the other operations in the network, including the Fourier layers, pointwise linear neural networks, and pointwise linear operators, are also implemented in a differentiable manner. This design eliminates issues related to subgradients or undefined derivatives at specific points and preserves the validity of higher-order derivative computations using automatic differentiation. A discussion of using Autograd with non-differentiable components is included in Appendix˜G.
4 Numerical Experiments
We use Meta-PDE (Qin et al., 2022) and the popular finite element method (FEM) FEniCS (Alnæs et al., 2015; Logg et al., 2011) as baselines. Meta-PDE learns initializations for PINNs over multiple instances that can be fine-tuned on any single instance, and two versions have been proposed based on the meta-learning algorithms MAML (Finn et al., 2017) and LEAP (Flennerhag et al., 2019). While FEniCS has been successfully applied in various disciplines, it has not been highly optimized for our applications and could possibly be outperformed by other FEMs and non-FEMs (Liu, 2009) such as Radial Basis Function (RBF) methods (interpolation using RBFs), Finite Point Methods (FPMs) and Moving Least Squares (MLS) methods (weighted least squares to approximate solutions) and spectral methods (expanding the solution in a basis of functions).
@@ 4.1 Burgers’ Equation
4.1.1 Problem Description
$u(x,0)$ $=u_{0}(x).$
We focus on the dataset of Li et al. (2021a; b) consisting of 800 instances of the Burgers’ equation with viscosity coefficient $\nu=0.01$ and $128\times 26$ resolution. Each sample in the dataset corresponds to a different initial condition $u_{0}$ drawn from the Gaussian process $\mathcal{N}(0,625(-\Delta+25I)^{-2})$. Examples of initial conditions and solutions to this time-dependent Burgers’ equation are shown in Figures˜4 and 5.
Figure 4: Examples of initial conditions used for the Burgers’ Equation.
$v=(\mathcal{G}_{m\mathrm{GNO}}\ \circ\ \mathcal{G}_{\mathrm{FNO}})(u_{0})$ (13)
4.1.2 Comparison of Differentiation Methods and Training Strategies
We trained models using $\mathcal{L}_{\text{Burgers}}$, where the derivatives are computed using various differentiation methods. The results, obtained by fine-tuning the models one instance at a time and then averaging over the dataset, are presented in Table˜1. Compared to numerical differentiation, $m\mathrm{GNO}\circ\mathrm{FNO}$ with Autograd was easier to tune and produced a model with data loss 20× lower (although 6× slower per training epoch), suggesting it better captures the physics underlying the data. It also achieved a physics loss 3000× smaller, but this was expected since the physics loss is evaluated using Autograd.
Table 1: Data and physics losses (computed using Autograd) for the proposed $m\mathrm{GNO}\circ\mathrm{FNO}$ models and PINO baselines, trained with $\mathcal{L}_{\text{Burgers}}$ where derivatives are computed using different methods.
PDE Residual Loss Relative L2 Data Loss
Autograd (ours) $1.62\cdot 10^{-6}$ $1.33\cdot 10^{-2}$
@@ 4.1.2 Comparison of Differentiation Methods and Training Strategies
Transolver Baseline (finite-differences) $3.01\cdot 10^{-2}$ $7.47\cdot 10^{-2}$
@@
A comparison to PINO (i.e. using an FNO instead of a $m\mathrm{GNO}\circ\mathrm{FNO}$) shows that PINO achieved a lower physics loss, but failed to reduce the Relative L2 data loss below $10\%$ on the hyperparameter sweep considered. Note that Autograd $m\mathrm{GNO}\circ\mathrm{FNO}$ was only 1.2× slower than Autograd PINO (although this was with a batch size of 1 and Autograd PINO could be accelerated more easily using batching). We also compared performance to Transolver (Wu et al., 2024) with finite differences. We performed a grid search over learning rate, hidden dimension, number of heads and layers, initial condition loss coefficient, and reported the best result in Table˜1. Transolver does not reduce the PDE residual as effectively as the other approaches, but strikes a better balance by achieving lower data loss than the remaining baselines. Nonetheless, it still falls short of the performance of Autograd $m\mathrm{GNO}\circ\mathrm{FNO}$, which attains a substantially lower PDE residual while staying closer to the reference data.
@@
We trained $m\mathrm{GNO}\circ\mathrm{FNO}$ models using a data loss, a physics loss, and a hybrid loss (data and physics). The results displayed in Table˜2 show that using the data loss helps reduce the gap with reference data, but can come at the cost of a higher physics loss, and vice versa.
@@
Table 2: Data and PDE loss of three models trained with different losses $\mathcal{L}$, using Autograd to compute the physics losses, for the Burgers’ equation.
$\mathcal{L}_{\text{data}}$ $\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}\ \ \$$\mathcal{L}_{\text{physics}}$
   Relative L2 Data Loss $1.42\cdot 10^{-3}$ $3.81\cdot 10^{-3}$ $1.33\cdot 10^{-2}$
   PDE Residual Loss $4.80\cdot 10^{-4}$ $1.48\cdot 10^{-4}$ $1.62\cdot 10^{-6}$
We also tried pre-training the $m\mathrm{GNO}\circ\mathrm{FNO}$ models using the data loss before fine-tuning them with the physics loss $\mathcal{L}_{\text{Burgers}}$ only, but were not able to obtain better results this way. The data loss rapidly deteriorates from its original value obtained in a data-driven way, while the physics loss improves slowly but does not get better than when training with hybrid and physics-only losses without data loss pre-training.
4.1.3 Subsampling and Data Efficiency
We now consider randomly subsampling the points at which the PDE residuals are evaluated to consider the effect of reducing the resolution used for the physics loss on performance. Table˜3 shows that despite the random locations and reduction of the number of points used to evaluate the derivatives, we maintain good results until the number of points becomes very low.
@@ 4.1.3 Subsampling and Data Efficiency
Table 3: PDE Residual and relative data losses for $m\mathrm{GNO}\circ\mathrm{FNO}$ models trained with the physics loss $\mathcal{L}_{\text{Burgers}}$ using Autograd derivatives evaluated at different numbers of randomly subsampled points from the original $128\times 26$ regular grid with 3328 points.
PDE Residual Loss Relative L2 Data Loss
@@
   Random 250 points $7\cdot 10^{-5} 1.441$$4\cdot 10^{-2}$
@@
We also investigate data efficiency of hybrid loss training, and more precisely how performance changes when varying the number of points at which the data loss and physics loss are evaluated when training the proposed $m\mathrm{GNO}\circ\mathrm{FNO}$ models using a hybrid training loss $\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}$. Table˜4 shows the results obtained with $m\mathrm{GNO}\circ\mathrm{FNO}$ models on the time-dependent Burgers’ equation. We see that performance remains unaffected until the spatial resolution falls below 16 or the temporal resolution falls below 5. This corresponds to 8× spatial subsampling and 5× time subsampling from the original $128\times 26$ grid, and highlights the data efficiency of the proposed approach: instead of requiring high resolution $128\times 26$ simulations (3328 points) for the data loss, we can obtain results of similar accuracy by supplementing low resolution simulations (e.g. at $22\times 9$, i.e. 198 points) with a physics loss evaluated at a higher resolution.
Table 4: Relative L2 data loss and physics loss (computed using Autograd) for the proposed $m\mathrm{GNO}\circ\mathrm{FNO}$ models on Burgers’ equation, when training with a hybrid loss. We consider the effect of reducing the resolution at which the data loss is evaluated during training, starting from the original $128\times 26$ resolution. Each row corresponds to a different spatial resolution (subsampled from 128), while each column corresponds to a different temporal resolution (subsampled from 26). We show in orange the settings where the data loss starts being affected, and in red the settings where the data loss is significantly worse.
Time Space $26$ $13$ 99 55 Physics $|$ Data Physics $|$ Data Physics $|$ Data Physics $|$ Data $128$ $1.48\cdot 10^{-4}\ |\ 3.81\cdot 10^{-3}$ $1.85\cdot 10^{-4}\ |\ 4.69\cdot 10^{-3}$ $77\cdot 10^{-4}\ |\ 4.80\cdot 10^{-3} 1.1$$11\cdot 10^{-4}\ |\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}6.84\cdot 10^{-3}}$  $64$ $1.73\cdot 10^{-4}\ |\ 3.83\cdot 10^{-3}$ $1.95\cdot 10^{-4}\ |\ 4.89\cdot 10^{-3}$ $86\cdot 10^{-4}\ |\ 4.82\cdot 10^{-3} 1.1$$17\cdot 10^{-4}\ |\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}7.04\cdot 10^{-3}}$  $43$ $1.74\cdot 10^{-4}\ |\ 4.04\cdot 10^{-3}$ $5\cdot 10^{-4}\ |\ 4.84\cdot 10^{-3} 1.881$$8\cdot 10^{-4}\ |\ 5.02\cdot 10^{-3}$ $1.14\cdot 10^{-4}\ |\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}6.86\cdot 10^{-3}}$  $32$ $1.74\cdot 10^{-4}\ |\ 4.22\cdot 10^{-3}$ $1.85\cdot 10^{-4}\ |\ 4.89\cdot 10^{-3}$ $82\cdot 10^{-4}\ |\ 4.91\cdot 10^{-3} 1.1$$17\cdot 10^{-4}\ |\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}6.84\cdot 10^{-3}}$  $26$ $1.82\cdot 10^{-4}\ |\ 4.21\cdot 10^{-3}$ $1.77\cdot 10^{-4}\ |\ 4.82\cdot 10^{-3}$ $92\cdot 10^{-4}\ |\ 4.87\cdot 10^{-3} 1.1$$13\cdot 10^{-4}\ |\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}6.88\cdot 10^{-3}}$  $22$ $1.70\cdot 10^{-4}\ |\ 4.03\cdot 10^{-3}$ $1.68\cdot 10^{-4}\ |\ 4.60\cdot 10^{-3}$ $1.72\cdot 10^{-4}\ |\ 4.77\cdot 10^{-3}$ $1.21\cdot 10^{-4}\ |\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}7.03\cdot 10^{-3}}$  $16$ $1.84\cdot 10^{-4}\ |\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}8.60\cdot 10^{-3}}$ $1.99\cdot 10^{-4}\ |\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}8.08\cdot 10^{-3}}$ $1.97\cdot 10^{-4}\ |\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}7.27\cdot 10^{-3}}$ $41\cdot 10^{-4}\ |\ {\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}8.33\cdot 10^{-3}}  11$11 $1.86\cdot 10^{-4}\ |\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.32\cdot 10^{-2}}$ $1.90\cdot 10^{-4}\ |\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.17\cdot 10^{-2}}$ $2.04\cdot 10^{-4}\ |\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.11\cdot 10^{-2}}$ $1.40\cdot 10^{-4}\ |\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.13\cdot 10^{-2}}$  77 $1.57\cdot 10^{-4}\ |\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}5.67\cdot 10^{-2}}$ $1.68\cdot 10^{-4}\ |\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}5.15\cdot 10^{-2}}$ $2.01\cdot 10^{-4}\ |\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}5.06\cdot 10^{-2}}$ $1.21\cdot 10^{-4}\ |\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}5.12\cdot 10^{-2}}$
4.1.4 Robustness to Noisy Data
We now investigate the effect of noisiness on reference data when training the proposed $m\mathrm{GNO}\circ\mathrm{FNO}$ models using a hybrid training loss $\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}$, both with finite differences and automatic differentiation.
More precisely, instead of the distance to the reference solution $u$, the data loss is based on the distance to a noisy reference solution
$\tilde{u}(x,t)=u(x,t)+\eta\cdot\varepsilon(x,t)\cdot\int_{0}^{1}\!\!\!\int_{0}^{1}\!|u(x^{\prime},t^{\prime})|\,\mathrm{d}x^{\prime}\,\mathrm{d}t^{\prime},$ (15)
where $\eta$ is the noise level, the integral represents the average absolute value of $u(x,t)$, and $\varepsilon(x,t)$ is an independent standard normal random variable for each $(x,t)$, that is $\varepsilon\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$.
Table˜5 illustrates the effect of noise in the reference data on the $m\mathrm{GNO}\circ\mathrm{FNO}$ models trained with a hybrid physics-data loss. The results show that the Autograd approach, our proposed method, maintains very low PDE residuals even as noise increases, demonstrating its robustness and stability. The relative L2 error with respect to the true solution grows with noise level, reflecting the unavoidable discrepancy between noisy training data and the true solution, but increases significantly slower than the training L2 error with respect to the noisy reference. This indicates that the physics loss truly helps the model to learn effectively the dynamics without overfitting to noise.
Compared to finite differences, Autograd achieves slightly lower residuals and errors across noise levels, highlighting its reliability for accurately computing derivatives within the physics loss. These results confirm that the proposed Autograd-based hybrid training is effective and resilient, preserving model accuracy even when the reference data is noisy, while finite differences provide a reasonable but less precise baseline.
Table 5: Effect of injecting noise of different levels $\eta$ (as defined in equation 15) in the reference data when training a $m\mathrm{GNO}\circ\mathrm{FNO}$ on Burgers’ equation with a hybrid loss. We report the PDE residual and the relative L2 data loss both to the true reference solution $u$ and to the noisy reference solution $\tilde{u}$. We consider both the cases where the derivatives are computed using Autograd and finite differences in the physics losses.
     Noise Level $\eta$    PDE Residual Relative L2 (true) Training Relative L2 (noisy)
Autograd
     $\eta=0\%$ (Reference) $3.43\cdot 10^{-4}$ $1.68\cdot 10^{-3}$ $1.68\cdot 10^{-3}$
     $\eta=0.1\%$ $3.39\cdot 10^{-4}$ $1.70\cdot 10^{-3}$ $1.90\cdot 10^{-3}$
     $\eta=0.5\%$ $3.08\cdot 10^{-4}$ $1.99\cdot 10^{-3}$ $4.62\cdot 10^{-3}$
     $\eta=1\%$ $2.79\cdot 10^{-4}$ $2.64\cdot 10^{-3}$ $8.74\cdot 10^{-3}$
     $\eta=2\%$ $3.01\cdot 10^{-4}$ $4.09\cdot 10^{-3}$ $1.72\cdot 10^{-2}$
     $\eta=5\%$ $4.93\cdot 10^{-4}$ $8.41\cdot 10^{-3}$ $4.29\cdot 10^{-2}$
     $\eta=10\%$ $5.73\cdot 10^{-4}$ $1.31\cdot 10^{-2}$ $8.58\cdot 10^{-2}$
     $\eta=20\%$ $9.79\cdot 10^{-4}$ $2.22\cdot 10^{-2}$ $1.70\cdot 10^{-1}$
     $\eta=50\%$ $1.67\cdot 10^{-3}$ $4.39\cdot 10^{-2}$ $3.97\cdot 10^{-1}$
Finite Differences
     $\eta=0\%$ (Reference) $4.78\cdot 10^{-4}$ $1.35\cdot 10^{-3}$ $1.35\cdot 10^{-3}$
     $\eta=0.1\%$ $4.88\cdot 10^{-4}$ $1.40\cdot 10^{-3}$ $1.63\cdot 10^{-3}$
     $\eta=0.5\%$ $5.00\cdot 10^{-4}$ $2.24\cdot 10^{-3}$ $4.73\cdot 10^{-3}$
     $\eta=1\%$ $5.30\cdot 10^{-4}$ $3.54\cdot 10^{-3}$ $9.10\cdot 10^{-3}$
     $\eta=2\%$ $6.28\cdot 10^{-4}$ $5.81\cdot 10^{-3}$ $1.78\cdot 10^{-2}$
     $\eta=5\%$ $8.23\cdot 10^{-4}$ $1.12\cdot 10^{-2}$ $4.40\cdot 10^{-2}$
     $\eta=10\%$ $1.16\cdot 10^{-3}$ $1.85\cdot 10^{-2}$ $8.74\cdot 10^{-2}$
     $\eta=20\%$ $1.71\cdot 10^{-3}$ $3.07\cdot 10^{-2}$ $1.72\cdot 10^{-1}$
     $\eta=50\%$ $2.98\cdot 10^{-3}$ $6.08\cdot 10^{-2}$ $4.01\cdot 10^{-1}$